Fill in the missing values in the function table:
Input(x) Output (y)
x=1 y=8
x=2 y=14
x=3 y=Unknown value
x=Unknown value y=24
x=Unknown value y=28
x=6 y=Unknown value
Please help me. I am stumped!

Answer by stanbon(75887)

Fill in the missing values in the function table:
Input(x) Output (y)
x=1 y=8
x=2 y=14
x=3 y=Unknown value
x=Unknown value y=24
x=Unknown value y=28
x=6 y=Unknown value
----------------------
Using the 1st two lines you have (1,8) and (2,14)
slope = (14-8)/(2-1) = 6
intercept: 8 = 6*1 + b
b = 2
--------
EQUATION:
y = 6x + 2
---------------
Using that equation you can fill in the missing values.
==========================
Cheers,
Stan H.